OpenStudy (anonymous):

g(n)= 4n-2 f(n)= n^2+4 Find g(-2) - f(-2)

OpenStudy (anonymous):

plug in -2 for the first and -2 for the second equation

OpenStudy (anonymous):

i want to compare my answer with yours

OpenStudy (anonymous):

what is your answer?

OpenStudy (anonymous):

i got -18

OpenStudy (anonymous):

g(-2)=4(-2)-2=-10, f(-2)=(-2)^2+4=8 g(-2)-f(-2)=-18

OpenStudy (anonymous):

-10 for the first

OpenStudy (anonymous):

and 8 for the second

OpenStudy (anonymous):

8 for the second

OpenStudy (anonymous):

so adding would be-2

OpenStudy (anonymous):

it says to subtract though..

OpenStudy (anonymous):

okay so -18

OpenStudy (anonymous):

lol thanks for the help!

OpenStudy (anonymous):

Your welcome :)